The vapor pressure, P, of a certain liquid was measured at two temperatures, T. If you were going to graphically determine the enthalpy of vaporizaton, AHvap, for this liquid, what points would you plot? The data is shown in this table Number Number T (K) P (kPa) 325 625 To avoid rounding errors, use three significant figures in the x values and four significant figures in the y values point 1:0.003 1.1694 3.22 7.03 Number Number Keep the pressure units in kPa point 2: 0.001 1.9502 Determine the rise, run, and slope of the line formed by these points rise run slope Number Number Number -0.002 0.7862 一0.0025 What is the enthalpy of vaporization of this liquid? Number J/ mol vap
Answer #1

X and Y coordinates will be decided by Clausius-Clapeyron equation

ln P =( \DeltaH/R) * (1/T)

y = m * x

1/T(K) = x

ln P(kPa) = y

x y

Point 1    (x1) 1/325= 0.00307K (y1) ln(3.22)= 1.1694 kPa
Point 2 (x2) 1/625=0.0016K (y2) ln(7.03)= 1.9502 kPa


Rise = y2-y1

= 1.9502 - 1.1694

= 0.7808

Run = x2-x1

= 0.0016 -0.00307

= - 0.00147

Slope= (y2-y1/x2-x1)

= 0.7808/(-0.00147)

= -531.156

Negative sign represents the downward slope

Hvap= m x R

= 531.156 x 8.314

= 4416.034